OpenStudy (anonymous):

College Algebra: reviewing for Final Exam! If f(x)=x^2+1 and g(x)=x-2, then (f of g)(x)=? I thought I was doing it correctly, but I didn't get the right answer.

OpenStudy (perl):

Did you get f (g(x)) = f ( x-2) = (x-2)^2 + 1

OpenStudy (anonymous):

Yes I did

OpenStudy (perl):

does it say you need to expand it

OpenStudy (anonymous):

I didn't get any more instructions. I thought the answer was x^2+5, but it's not. This is how I did it.